OpenStudy (anonymous):

5. A 40% salt solution is to be mixed with a 60% salt solution to obtain 100 gallons of a 45% salt solution. How much of each solution must be used. 6. A field is twice as long as wide. The length of the diagonal is 64 feet. What are the dimensions of the field? (Round answer to nearest tenth.) 8. Two trains leave at the same time from the same station, one traveling due south and the other traveling due east. The eastbound train is traveling 20 miles an hour faster than the southbound train. After 4 hours the trains are 400 miles apart, what is the rate of each train.

OpenStudy (anonymous):

can anyone answer any of these questions? PLEASE!

OpenStudy (anonymous):

Second One: A field is twice as long as it is wide. Let its width be : x Then Length will be : 2x \[(Diagonal)^2 = (Length)^2 + (Width)^2\] \[x^2 + 4x^2 = 64^2\] \[5x^2 = 64^2\] Taking square root both the sides, \[x \sqrt{5} = 64\] \[x = \frac{64}{2.236}\] Divide to get the answer..

OpenStudy (anonymous):

thank u

OpenStudy (anonymous):

Welcome dear..

OpenStudy (anonymous):

can anyone figure this one out? One positive number is one-fifth of another number. The difference between the two numbers is 76. Find the numbers.

OpenStudy (anonymous):

or any of the other ones

hero (hero):

Hint for no 5: x + y = 100 .4x + .6y = 45

OpenStudy (anonymous):

Let the other number is : x So, What is the value of One Positive Number?? It is : x/5 So, x - x/5 = 76 Now solve for x..

OpenStudy (anonymous):

@Hero I think it is a question of mixing and not linear equations..

OpenStudy (cwrw238):

let x be volume of 40% and y the volume of 60 % solution x + y = 100 0.40x + 0.60y = 45 solve for x and y

hero (hero):

waterineyes, you've never solved such equations before?

OpenStudy (anonymous):

No, but it is looking like if they are not from Linear Equations..

hero (hero):

You mean systems of equations. That's exactly what it is

OpenStudy (anonymous):

you guys are life savers do u think you could solve these? A plane is flying nonstop from Pittsburgh to Los Angles, a distance of about 2400 miles. After 1 hour 15 minutes in the air, the plane flies over St Louis (a distance of 800 miles from Pittsburgh). Estimate the time the plane will arrive in Los Angles if it departs from Pittsburgh at 12 p.m.(noon). You are driving on an interstate highway to a town that is 420 miles from your home. After 2 hours driving, you pass a town that you know is 120 miles from your home. Assuming that you continue at the same rate, except for a half hour food break, how long will it take for you to get to your destination?

OpenStudy (ganpat):

for the plane one, speed = distance / time pitesburgh to St Louis ; speed = 800/1.25 = 640 mph.. Therefore from pitsbugh to LA = speed = distance / time 640 = 2400/ time time = 2400/640 = 3.75... i.e. 3 hrs and 45 minutes from Pitsburgh to LA.. time of arrival 15:45...

OpenStudy (ganpat):

try using same step for second one as well...